OpenStudy (fanduekisses):

What is the exact value of 20z over sqrt z^13 ?

OpenStudy (anonymous):

do you mean \[20z \div z ^{13}\]

OpenStudy (fanduekisses):

No, I mean \[\left(\begin{matrix}20z \\ \sqrt{z ^{13}}\end{matrix}\right)\]